# Adaptive finite element interpolated neural networks

### v1

@@ 3.2. Quasi-emulation
In the previous section, we have considered the case in which the NN space cannot emulate the FE space, i.e., $\mathcal{N}_{h}\subsetneq U_{h}$. If the NN is expressive enough, we can prove that $\mathcal{N}_{h}=U_{h}$ following the construction in [17, Prop. 3.2].
@@
The proof follows the same lines as in [17, Prop. 3.2]. The neural network above can represent in the last layer all the shape functions in the adaptive FE space $U_{h}$, both the free ones as the ones that are constrained and associated to hanging nodes. The linear combination of these last layer shape functions can represent any function in $U_{h}$. As a result, the interpolation operator $\pi_{h}$ is surjective. ∎
@@
The previous result is overly pessimistic and can be improved when using meshes that are obtained after adaptive mesh refinement of a regular mesh. In this case, the FE space $U_{h}$ can be represented by a NN with a much smaller number of neurons, since one can exploit a tensor product structure in the proof (see [17]).
@@ 3.3. Generalisation error and equivalence classes
In order to better bound the total error of the NN, we can make use of quotient spaces (see a related approach in [18] for VPINNs). Let us define the following equivalence relation $\sim$: two NN realisations $\mathcal{N}(\boldsymbol{\theta})$ and $\mathcal{N}(\boldsymbol{\theta}^{\prime})$ are equivalent if their interpolations on the adaptive FE space $U_{h}$ are identical, i.e., $\pi_{h}(\mathcal{N}(\boldsymbol{\theta}))=\pi_{h}(\mathcal{N}(\boldsymbol{\theta}^{\prime}))$. The equivalence class of $\mathcal{N}(\boldsymbol{\theta})$ is denoted by $[\mathcal{N}(\boldsymbol{\theta})]$. We can define the quotient space $\mathcal{N}/\sim$ as the set of all equivalence classes $[\mathcal{N}(\boldsymbol{\theta})]$.
@@
The first term in the error estimate is the FE error, the second term is the generalisation error, and the third term is a training error. The generalisation error describes the maximum distance between the FE solution and the NN realisations in the corresponding equivalence class. Since the loss function does not feel $\|\pi_{h}^{\perp}(u_{\mathcal{N}})\|$, there is no explicit control on it. This error is related to the expressiveness of the NN with respect to the FE mesh. The more expressive the NN, the larger the generalisation error can be, as it is well-known in the machine learning literature (over-fitting). Besides, if the NN is too coarse, the emulation assumption will not hold and the FEINN will not be able to capture the FE solution. 111 This discussion motivates the usage of a regularisation term that controls the distance between the NN and the FE solution. For example, one could consider a term of the order $h^{-2}\|\pi^{\perp}(u_{\mathcal{N}})\|_{L^{2}(\Omega)}$ or $\|\pi^{\perp}(u_{\mathcal{N}})\|_{H^{1}(\Omega)}$, integrated with a high order quadrature. Such a term is weakly consistent, i.e., it does not affect the convergence of FEINNs compared to FEM. This stabilisation resembles the so-called orthogonal subscale stabilisation common in the FE literature to circumvent the inf-sup condition (see, e.g., [37, 38]). We have not considered stabilisation in this work.
The previous analysis is FE-centric, in the sense that we consider the NN error by comparing it with the FE solution and use FE theory to quantify the error. Even though the FEINN can capture the FE solution with a much smaller number of parameters than the number of DoFs in the FE space [17], it can be argued that if the only aim is to attain the FE error, it is more effective to solve the FE problem. However, the FEINN approach can naturally deal with inverse problems by simply adding data misfit terms to the loss function and the NN parametrisation regularises the inverse problem and can readily be used in a one-shot fashion, avoiding the need of adjoint problems [17]. It has also been observed experimentally in [17] that the FE-centric analysis of FEINNs is overly pessimistic in general. The trained NN using the FEINN approach can be up to two orders of magnitude more accurate than the FE solution for forward problems. In Sec. 5, we analyse this fact in the case of singular solution and sharp gradients combined with $h$-adaptivity and a posteriori error estimation. A mathematical justification of the superior accuracy of the FEINN trained NN compared to the FEM solution is still open.
@@ 4.1. Automatic mesh adaptation using forest-of-octrees
Automatic mesh adaptation is a key component of $h$-adaptive FEINNs. In this work, we leverage a particular kind of hierarchically-adapted meshes known as forest-of-octrees (see, e.g., [39, 40]). Forest-of-octrees meshes can be seen as a two-level decomposition of $\Omega$, referred to as macro and micro level, respectively. The macro level is a suitable conforming partition $\mathcal{C}$ of $\Omega$ into quadrilateral ($d=2$) or hexahedral cells ($d=3$). This mesh, which may be generated, e.g., using an unstructured mesh generator, is referred to as the coarse mesh. At the micro level, each of the cells of $\mathcal{C}$ becomes the root of an adaptive octree with cells that can be recursively and dynamically refined or coarsened using the so-called $1:2^{d}$ uniform partition rule. If a cell is marked for refinement, then it is split into $2^{d}$ children cells by subdividing all parent cell edges. If all children cells of a parent cell are marked for coarsening, then they are collapsed into the parent cell. The union of all leaf cells in this hierarchy forms the decomposition of the domain at the micro level, i.e., $\mathcal{T}$.
In this work, we utilise the GridapP4est.jl [41] Julia package in order to handle such kind of meshes. This package, built upon the p4est meshing engine [39], is endowed with the so-called Morton space-filling curves, and it provides high-performance and low-memory footprint algorithms to handle forest-of-octrees.
@@ 4.2. Non-conforming interpolation of neural networks
Unlike our previous work [17], which relies on (uniformly-refined) conforming meshes, in this work we interpolate the NN onto (grad-conforming) FE spaces built out of non-conforming forest-of-octrees meshes. With this kind of meshes, generating a cell-conforming global enumeration of DoFs is not sufficient to impose the required continuity at the interface of cells with different refinement levels. In order to remedy this, it is standard practice to equip the FE space with additional linear multi-point constraints at the DoFs lying on hanging vertices, edges and faces. The structure and set up of such constraints is well-established knowledge, and thus not covered here; see, e.g., [40, 42] for further details. In any case, the presence of these linear multi-point constraints in the FE space implies that the interpolation of the NN at hanging DoFs cannot by given by its evaluation at the corresponding nodes, but rather by a linear combination of the NN evaluations at the nodes of the adjacent coarser cell. To this end, GridapP4est.jl provides the required functionality to build and apply these linear multi-point constraints, and to resolve these constraints during FE assembly.
@@ 4.3. Finite element interpolated neural networks
In this section we briefly overview the main building blocks of our Julia implementation of $h$-adaptive FEINNs. For a more complete coverage of these (including the inverse problem case), we refer to [17]. At a purely algebraic level, we can rewrite the preconditioned loss (12) as:
@@
where $r_{h}\in V_{h}$ is the discrete Riesz projection of the FE residual. The vector of DoFs of $r_{h}$ is $\mathbf{B}^{-1}(\mathbf{A}\mathbf{u}_{h}-\mathbf{f})$, where $\mathbf{B}$ is the discretisation in the test space $V_{h}$ of the Gram matrix associated to the inner product in $U$, e.g., the Laplacian matrix for the Poisson problem, $\mathbf{A}$ is the coefficient matrix resulting from discretisation, $\boldsymbol{\theta}$ is the parameters for the NN $u_{\mathcal{N}}$, $\mathbf{u}_{h}$ is the vector of DoFs of $U_{h}$, and $\mathbf{f}$ is the right-hand side (RHS) vector. We use the abstract data structures in Gridap.jl [43, 44] as implemented by GridapP4est.jl to compute $\mathbf{A}$, $\mathbf{f}$, and the FE nodal coordinates $\mathbf{X}$ of those DoFs which are not subject to linear multi-point constraints. We utilise Flux.jl [45, 46] to build $u_{\mathcal{N}}$ and evaluate $u_{\mathcal{N}}$ at $\mathbf{X}$ to compute $\mathbf{u}_{h}$. Without preconditioning, i.e., $\mathbf{B}=\mathbf{I}$, we can minimise the loss $\left\lVert\mathbf{A}\mathbf{u}_{h}(\boldsymbol{\theta})-\mathbf{f}\right\rVert_{\chi}$ directly. During training, we rely on Zygote.jl [47] to compute the gradient of the loss with respect to $\boldsymbol{\theta}$ (see [17] for details). We note that in the context of $h$-adaptive FEINNs, we need to recompute $\mathbf{A}$, $\mathbf{f}$, and $\mathbf{u}_{h}$ at each adaptation step. Furthermore, depending on the specific problem, these arrays can be stored cell-wise to improve efficiency.
@@
The forward and backward passes of the above steps are both handled by Gridap.jl. It is noteworthy that in step (1), we do not explicitly compute the inverse of $\mathbf{B}$, but rather solve the linear system $\mathbf{B}\mathbf{r}_{h}=\mathbf{A}\mathbf{u}_{h}-\mathbf{f}$ to compute the vector of DoFs $\mathbf{r}_{h}$ that uniquely determine $r_{h}\in V_{h}$. Considering the problem linear, we can reuse the factorisation of $\mathbf{B}$, leading to linear complexity computational cost in the application of $\mathbf{B}^{-1}$ per iteration. Alternatively, as explored in [17], we can use a cheaper preconditioner $\mathbf{B}$, such as the GMG preconditioner, which has linear complexity in the number of DoFs of the FE space. Besides, given that $r_{h}\in V_{h}$ and $V_{h}$ is a linearised FE space, and the gradient of $r_{h}$ is piecewise constant, the computation of $\left\lVert r_{h}\right\rVert_{X}$ is very cheap.
@@ 5. Numerical experiments
Tab. 1 lists the NN architectures and training iterations for the different problems tackled in this section. For 2D experiments, we adopt the NN architecture in [17], namely $L=5$ layers and $n=50$ neurons for each hidden layer. For 3D experiments, since we work with a more complex problem with more DoFs, we increase the number of layers to $L=10$. In addition, we use $\rho=\tanh$ as the activation function, the Glorot uniform method [48] as the NN parameter initialisation strategy. We employ the BFGS optimiser in Optim.jl [49] to train the NNs, as recommended in [17]. Moreover, in our preliminary experiments, we have observed that the results are generally robust to variations in NN initialisations, provided that the initial mesh is sufficiently fine and the NN is adequately trained. As a result, at each adaptation step, we heuristically set the number of the training iterations as a function of the number of DoFs in the trial FE space. In particular, the higher the number of DoFs, the higher the number of training iterations. We acknowledge that advanced/automatic optimiser termination conditions could result in better accuracy/performance trade-offs than the ones shown in this section. However, developing such conditions is beyond the scope of this paper.
@@
The unpreconditioned loss (13) has proven effective in various cases, as shown in [17]. Therefore, we use it as the default loss function in the experiments. The preconditioned loss function (12) is only used in Sec. 5.1.2 to examine the effectiveness of the proposed preconditioning strategy. With regards to the norm choice in the loss (13), we use the $\ell^{1}$ norm. We note that other norms can also be explored. For example, our previous work [17] shows that both $\ell^{2}$ and $\ell^{1}$ norms in (13) produce similar results on (quasi-)uniform meshes. Besides, we explore different norm options in the preconditioned loss (12), as detailed in Sec. 5.1.2.
The main motivation of this work is to showcase the capability of the proposed method in handling PDEs with sharp gradients or singularities. For this purpose, we have considered standard tests on simple geometries. However, as demonstrated in our previous work [17], the FEINN method can also be applied to problems posed on more complex domains, and $h$-adaptive FEINNs inherit this capability seamlessly. To attack this kind of multi-scale problems, one can generate a forest-of-octrees built out of a coarse mesh $\mathcal{C}$ generated by an unstructured mesh generator such as, e.g., Gmsh [50]; see Sec. 4.1.
@@ 5.1.1. The comparison of error indicators
Regarding the errors in NNs themselves, as shown in the second column of Fig. 2, we observe that the majority of the $L^{2}$ errors from the 20 experiments are within an acceptable range. This is evident as all the lines (medians) corresponding to different error indicators closely align with those of the interpolated NNs. Nevertheless, the bands for $L^{2}$ and $H^{1}$ errors become noticeably wider, suggesting larger variance in the results. The network error indicator stands out for its robustness, as evidenced by the narrowest bands in Fig. 2b and  2d. In our previous work [17], we have highlighted that for a smooth true solution, NNs have the potential to outperform the FEM solution on the same mesh. We observe a similar phenomenon here concerning $H^{1}$ errors: with real error indicator, NNs without interpolation can beat FEM solution (represented by the red dashed line) by approximately 2 orders of magnitude at the last adaptation step. For the remaining two error indicators, the majority of the $H^{1}$ errors fall below the FEINN line, indicating improved performance of the NNs compared to their interpolation.

### v2

@@ 3.2. Quasi-emulation
In the previous section, we have considered the case in which the NN space cannot emulate the FE space, i.e., $\mathcal{N}_{h}\subsetneq U_{h}$. If the NN is expressive enough, we can prove that $\mathcal{N}_{h}=U_{h}$ following the construction in [Badia2024, Prop. 3.2].
@@
The proof follows the same lines as in [Badia2024, Prop. 3.2]. The neural network above can represent in the last layer all the shape functions in the adaptive FE space $U_{h}$, both the free ones as the ones that are constrained and associated to hanging nodes. The linear combination of these last layer shape functions can represent any function in $U_{h}$. As a result, the interpolation operator $\pi_{h}$ is surjective. ∎
@@
The previous result is overly pessimistic and can be improved when using meshes that are obtained after adaptive mesh refinement of a regular mesh. In this case, the FE space $U_{h}$ can be represented by a NN with a much smaller number of neurons, since one can exploit a tensor product structure in the proof (see [Badia2024]).
@@ 3.3. Generalisation error and equivalence classes
In order to better bound the total error of the NN, we can make use of quotient spaces (see a related approach in [Rojas2023] for VPINNs). Let us define the following equivalence relation $\sim$: two NN realisations $\mathcal{N}(\boldsymbol{\theta})$ and $\mathcal{N}(\boldsymbol{\theta}^{\prime})$ are equivalent if their interpolations on the adaptive FE space $U_{h}$ are identical, i.e., $\pi_{h}(\mathcal{N}(\boldsymbol{\theta}))=\pi_{h}(\mathcal{N}(\boldsymbol{\theta}^{\prime}))$. The equivalence class of $\mathcal{N}(\boldsymbol{\theta})$ is denoted by $[\mathcal{N}(\boldsymbol{\theta})]$. We can define the quotient space $\mathcal{N}/\sim$ as the set of all equivalence classes $[\mathcal{N}(\boldsymbol{\theta})]$.
@@
The first term in the error estimate is the FE error, the second term is the generalisation error, and the third term is a training error. The generalisation error describes the maximum distance between the FE solution and the NN realisations in the corresponding equivalence class. Since the loss function does not feel $\|\pi_{h}^{\perp}(u_{\mathcal{N}})\|$, there is no explicit control on it. This error is related to the expressiveness of the NN with respect to the FE mesh. The more expressive the NN, the larger the generalisation error can be, as it is well-known in the machine learning literature (over-fitting). Besides, if the NN is too coarse, the emulation assumption will not hold and the FEINN will not be able to capture the FE solution. 111 This discussion motivates the usage of a regularisation term that controls the distance between the NN and the FE solution. For example, one could consider a term of the order $h^{-2}\|\pi^{\perp}(u_{\mathcal{N}})\|_{L^{2}(\Omega)}$ or $\|\pi^{\perp}(u_{\mathcal{N}})\|_{H^{1}(\Omega)}$, integrated with a high order quadrature. Such a term is weakly consistent, i.e., it does not affect the convergence of FEINNs compared to FEM. This stabilisation resembles the so-called orthogonal subscale stabilisation common in the FE literature to circumvent the inf-sup condition (see, e.g., [Codina2017, Badia2012]). We have not considered stabilisation in this work.
The previous analysis is FE-centric, in the sense that we consider the NN error by comparing it with the FE solution and use FE theory to quantify the error. Even though the FEINN can capture the FE solution with a much smaller number of parameters than the number of DoFs in the FE space [Badia2024], it can be argued that if the only aim is to attain the FE error, it is more effective to solve the FE problem. However, the FEINN approach can naturally deal with inverse problems by simply adding data misfit terms to the loss function and the NN parametrisation regularises the inverse problem and can readily be used in a one-shot fashion, avoiding the need of adjoint problems [Badia2024]. It has also been observed experimentally in [Badia2024] that the FE-centric analysis of FEINNs is overly pessimistic in general. The trained NN using the FEINN approach can be up to two orders of magnitude more accurate than the FE solution for forward problems. In Sec. 5, we analyse this fact in the case of singular solution and sharp gradients combined with $h$-adaptivity and a posteriori error estimation. A mathematical justification of the superior accuracy of the FEINN trained NN compared to the FEM solution is still open.
@@ 4.1. Automatic mesh adaptation using forest-of-octrees
Automatic mesh adaptation is a key component of $h$-adaptive FEINNs. In this work, we leverage a particular kind of hierarchically-adapted meshes known as forest-of-octrees (see, e.g., [Burstedde2011, Badia2020]). Forest-of-octrees meshes can be seen as a two-level decomposition of $\Omega$, referred to as macro and micro level, respectively. The macro level is a suitable conforming partition $\mathcal{C}$ of $\Omega$ into quadrilateral ($d=2$) or hexahedral cells ($d=3$). This mesh, which may be generated, e.g., using an unstructured mesh generator, is referred to as the coarse mesh. At the micro level, each of the cells of $\mathcal{C}$ becomes the root of an adaptive octree with cells that can be recursively and dynamically refined or coarsened using the so-called $1:2^{d}$ uniform partition rule. If a cell is marked for refinement, then it is split into $2^{d}$ children cells by subdividing all parent cell edges. If all children cells of a parent cell are marked for coarsening, then they are collapsed into the parent cell. The union of all leaf cells in this hierarchy forms the decomposition of the domain at the micro level, i.e., $\mathcal{T}$.
In this work, we utilise the GridapP4est.jl [GridapP4est2024] Julia package in order to handle such kind of meshes. This package, built upon the p4est meshing engine [Burstedde2011], is endowed with the so-called Morton space-filling curves, and it provides high-performance and low-memory footprint algorithms to handle forest-of-octrees.
@@ 4.2. Non-conforming interpolation of neural networks
Unlike our previous work [Badia2024], which relies on (uniformly-refined) conforming meshes, in this work we interpolate the NN onto (grad-conforming) FE spaces built out of non-conforming forest-of-octrees meshes. With this kind of meshes, generating a cell-conforming global enumeration of DoFs is not sufficient to impose the required continuity at the interface of cells with different refinement levels. In order to remedy this, it is standard practice to equip the FE space with additional linear multi-point constraints at the DoFs lying on hanging vertices, edges and faces. The structure and set up of such constraints is well-established knowledge, and thus not covered here; see, e.g., [Badia2020, Olm2019] for further details. In any case, the presence of these linear multi-point constraints in the FE space implies that the interpolation of the NN at hanging DoFs cannot by given by its evaluation at the corresponding nodes, but rather by a linear combination of the NN evaluations at the nodes of the adjacent coarser cell. To this end, GridapP4est.jl provides the required functionality to build and apply these linear multi-point constraints, and to resolve these constraints during FE assembly.
@@ 4.3. Finite element interpolated neural networks
In this section we briefly overview the main building blocks of our Julia implementation of $h$-adaptive FEINNs. For a more complete coverage of these (including the inverse problem case), we refer to [Badia2024]. At a purely algebraic level, we can rewrite the preconditioned loss (12) as:
@@
where $r_{h}\in V_{h}$ is the discrete Riesz projection of the FE residual. The vector of DoFs of $r_{h}$ is $\mathbf{B}^{-1}(\mathbf{A}\mathbf{u}_{h}-\mathbf{f})$, where $\mathbf{B}$ is the discretisation in the test space $V_{h}$ of the Gram matrix associated to the inner product in $U$, e.g., the Laplacian matrix for the Poisson problem, $\mathbf{A}$ is the coefficient matrix resulting from discretisation, $\boldsymbol{\theta}$ is the parameters for the NN $u_{\mathcal{N}}$, $\mathbf{u}_{h}$ is the vector of DoFs of $U_{h}$, and $\mathbf{f}$ is the right-hand side (RHS) vector. We use the abstract data structures in Gridap.jl [Gridap2020, Gridap2022] as implemented by GridapP4est.jl to compute $\mathbf{A}$, $\mathbf{f}$, and the FE nodal coordinates $\mathbf{X}$ of those DoFs which are not subject to linear multi-point constraints. We utilise Flux.jl [Flux2018, FluxOSS2018] to build $u_{\mathcal{N}}$ and evaluate $u_{\mathcal{N}}$ at $\mathbf{X}$ to compute $\mathbf{u}_{h}$. Without preconditioning, i.e., $\mathbf{B}=\mathbf{I}$, we can minimise the loss $\left\lVert\mathbf{A}\mathbf{u}_{h}(\boldsymbol{\theta})-\mathbf{f}\right\rVert_{\chi}$ directly. During training, we rely on Zygote.jl [Zygote2018] to compute the gradient of the loss with respect to $\boldsymbol{\theta}$ (see [Badia2024] for details). We note that in the context of $h$-adaptive FEINNs, we need to recompute $\mathbf{A}$, $\mathbf{f}$, and $\mathbf{u}_{h}$ at each adaptation step. Furthermore, depending on the specific problem, these arrays can be stored cell-wise to improve efficiency.
@@
The forward and backward passes of the above steps are both handled by Gridap.jl. It is noteworthy that in step (1), we do not explicitly compute the inverse of $\mathbf{B}$, but rather solve the linear system $\mathbf{B}\mathbf{r}_{h}=\mathbf{A}\mathbf{u}_{h}-\mathbf{f}$ to compute the vector of DoFs $\mathbf{r}_{h}$ that uniquely determine $r_{h}\in V_{h}$. Considering the problem linear, we can reuse the factorisation of $\mathbf{B}$, leading to linear complexity computational cost in the application of $\mathbf{B}^{-1}$ per iteration. Alternatively, as explored in [Badia2024], we can use a cheaper preconditioner $\mathbf{B}$, such as the GMG preconditioner, which has linear complexity in the number of DoFs of the FE space. Besides, given that $r_{h}\in V_{h}$ and $V_{h}$ is a linearised FE space, and the gradient of $r_{h}$ is piecewise constant, the computation of $\left\lVert r_{h}\right\rVert_{X}$ is very cheap.
@@ 5. Numerical experiments
Tab. 1 lists the NN architectures and training iterations for the different problems tackled in this section. For 2D experiments, we adopt the NN architecture in [Badia2024], namely $L=5$ layers and $n=50$ neurons for each hidden layer. For 3D experiments, since we work with a more complex problem with more DoFs, we increase the number of layers to $L=10$. In addition, we use $\rho=\tanh$ as the activation function, the Glorot uniform method [Glorot2010] as the NN parameter initialisation strategy. We employ the BFGS optimiser in Optim.jl [Optimjl2018] to train the NNs, as recommended in [Badia2024]. Moreover, in our preliminary experiments, we have observed that the results are generally robust to variations in NN initialisations, provided that the initial mesh is sufficiently fine and the NN is adequately trained. As a result, at each adaptation step, we heuristically set the number of the training iterations as a function of the number of DoFs in the trial FE space. In particular, the higher the number of DoFs, the higher the number of training iterations. We acknowledge that advanced/automatic optimiser termination conditions could result in better accuracy/performance trade-offs than the ones shown in this section. However, developing such conditions is beyond the scope of this paper.
@@
The unpreconditioned loss (13) has proven effective in various cases, as shown in [Badia2024]. Therefore, we use it as the default loss function in the experiments. The preconditioned loss function (12) is only used in Sec. 5.1.2 to examine the effectiveness of the proposed preconditioning strategy. With regards to the norm choice in the loss (13), we use the $\ell^{1}$ norm. We note that other norms can also be explored. For example, our previous work [Badia2024] shows that both $\ell^{2}$ and $\ell^{1}$ norms in (13) produce similar results on (quasi-)uniform meshes. Besides, we explore different norm options in the preconditioned loss (12), as detailed in Sec. 5.1.2.
The main motivation of this work is to showcase the capability of the proposed method in handling PDEs with sharp gradients or singularities. For this purpose, we have considered standard tests on simple geometries. However, as demonstrated in our previous work [Badia2024], the FEINN method can also be applied to problems posed on more complex domains, and $h$-adaptive FEINNs inherit this capability seamlessly. To attack this kind of multi-scale problems, one can generate a forest-of-octrees built out of a coarse mesh $\mathcal{C}$ generated by an unstructured mesh generator such as, e.g., Gmsh [Gmsh2009]; see Sec. 4.1.
@@ 5.1.1. The comparison of error indicators
Regarding the errors in NNs themselves, as shown in the second column of Fig. 2, we observe that the majority of the $L^{2}$ errors from the 20 experiments are within an acceptable range. This is evident as all the lines (medians) corresponding to different error indicators closely align with those of the interpolated NNs. Nevertheless, the bands for $L^{2}$ and $H^{1}$ errors become noticeably wider, suggesting larger variance in the results. The network error indicator stands out for its robustness, as evidenced by the narrowest bands in Fig. 2b and  2d. In our previous work [Badia2024], we have highlighted that for a smooth true solution, NNs have the potential to outperform the FEM solution on the same mesh. We observe a similar phenomenon here concerning $H^{1}$ errors: with real error indicator, NNs without interpolation can beat FEM solution (represented by the red dashed line) by approximately 2 orders of magnitude at the last adaptation step. For the remaining two error indicators, the majority of the $H^{1}$ errors fall below the FEINN line, indicating improved performance of the NNs compared to their interpolation.